OpenStudy (anonymous):

MULTIPLE CHOICE, PLEASE ANSWER What are the values of the variable in 37 = 30 − x2? A.−+i B.−+3i2√ C.−+2i3√ D.−+i7√

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@superdavesuper

OpenStudy (anonymous):

Note that \(\large 37=30-x^2 \implies x^2=-7\). What do you get when you square root both sides? :-)

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

It's not A. Let's think about a slightly different problem. Suppose you want to find the solution to \(\large x^2=5\). If you take square roots of both sides, we get \(\large x=\pm\sqrt{5}\), which can't be simplified any further. Now, going back to your equation \(\large x^2=-7\), what do you get when you square root both sides? Also keep in mind that since you'll be taking a square root of a negative number, there will be an \(\large i\) term in your answer. Does this clarify things? :-)

OpenStudy (anonymous):

Hmmm... D? :)

OpenStudy (anonymous):

Yes, D (\(\large x=\pm\sqrt{7}i\)) is the correct choice. Does that make a little more sense now? :-)